### v1

# Gradient-Mask Tuning Elevates the Upper Limits of LLM Performance

Abstract
Large language models (LLMs) have revolutionized lots of fields of research. Although it is well-known that fine-tuning is essential for enhancing the capabilities of LLMs, existing research suggests that there is potential redundancy in the fine-tuning process and therefore proposes to update only a subset of parameters. However, these methods fail to leverage the task-specific information to identify important parameters during training. Based on the insight that gradients inherently contain information on task-specific data, we propose Gradient-Mask Tuning (GMT), a method that selectively updates parameters during training based on their gradient information. Specifically, we compute the absolute values of the gradients and apply masking to those with relatively smaller magnitudes. Our empirical results across various tasks demonstrate that GMT not only outperforms traditional fine-tuning methods but also elevates the upper limits of LLM performance. Further analysis indicates that GMT exhibits insensitivity to mask ratio and possesses computational efficiency comparable to vanilla SFT.
Gradient-Mask Tuning Elevates the Upper Limits of LLM Performance
Haoling Li1,2thanks:   Work done during his internship at Microsoft., Xin Zhang 2thanks:   Corresponding author., Xiao Liu2, Yeyun Gong 2, Yifan Wang1, Yujiu Yang 1222https://evalplus.github.io/leaderboard.html, Qi Chen2, Peng Cheng 2 1Tsinghua University, 2Microsoft Research li-hl23@mails.tsinghua.edu.cnxinzhang3@microsoft.com
1 Introduction
Large language models (LLMs) have pivoted the centerpiece of various tasks such as textual understanding, natural language planning and instruction following (Achiam et al., 2023; Chowdhery et al., 2023; Touvron et al., 2023; Jiang et al., 2023). Although it is well-known that fine-tuning is essential to enhance the ability of LLMs on downstream tasks, existing works (Liu et al., 2024a; Yadav et al., 2024; Wang et al., 2024a) have demonstrated that potential redundancy will be leveraged during fine-tuning, suggesting that not all parameters need to be updated for LLMs.
Considering the above concerns, delta tuning (Ding et al., 2023), a fine-tuning framework that diminishes redundancy by updating only a portion of the model parameters, has been proposed. Compared to vanilla fine-tuning, existing delta tuning methods can be classified into two main categories based on implementation. The first is to randomly reset some of the parameters in the fine-tuned model checkpoints to those in the original model checkpoints after fine-tuning (Yu et al., 2023a), which is also called one-off drop. The second is to selectively update only a subset of the whole parameters during the model training process based on either random selection (Lee et al., 2019; Hui et al., 2024) or predefined rules (Zaken et al., 2022; Hu et al., 2023, 2021; Fu et al., 2023). Although these methods show the potential of achieving better results than vanilla fine-tuning and involve fewer changes to the original parameters, these is a problem that cannot be ignored that these methods tune the fixed set of parameters without leveraging information from task-specific data, leading to sub-optimal optimization target.
To mitigate this problem, an appropriate parameter selection method during training that can fully exploit the information of task-specific data is desired. Since the fact that the delta parameters (i.e., the disparity between fine-tuned and pre-trained parameters) are the overall accumulation of gradients during training, it is intuitive that using the magnitude of gradient will lead to an finer-grained dropping effect. In this work, we present the Gradient-Mask Tuning (GMT) to select the set of parameters to be updated during training based on the gradient information associated with the task-specific data. In a nutshell, after calculating the gradients in the backward process, we measure the absolute values of gradients and mask those with small values. The reasons we choose the magnanimity of gradient are two folds. First, gradient naturally conveys the information of task-specific data and is a by-product of without additional computations during training. Second, we believe that different network parameters vary in their importance for different training data. The gradient is a well-suited criterion for determining the importance of these parameters, which leads to finer-grained control of the entire fine-tuning process.
To evaluate the effectiveness of GMT, we provide theoretical analysis and conduct experiments with various base models on several widely-used benchmarks on math reasoning, code generation, and general instruction following. The experimental results demonstrate that GMT exploits the gradient information to identify more important network parameters for sparse updating with acceptable extra time spent, which is better compared to several representative baselines including vanilla fine-tuning, one-off drop, random mask and so on. Furthermore, it illustrates that GMT is not sensitive to mask ratio selection and is more effective than random parameter updating. We further analyze the computational FLOPs and time efficiency of GMT, demonstrating that it enables a plug-and-play replacement of vanilla SFT without destroying the network architecture.
Our contributions can be summarized as:
  • We propose GMT, which masks the gradients with small absolute values and discriminates the importance of parameters during training, naturally utilizing the information of task-specific data.
  • We conduct both theoretical analysis and exhaustive experiments with different base models on several benchmarks comparing with representative baselines. The results confirm the effectiveness of GMT.
  • We further demonstrate the adaptability, robustness and efficiency of the GMT by analyzing the drop strategy, mask ratio and time efficiency.
2 Methodology
In this section, we extend the description of delta tuning to facilitate further derivation and then detail the mathematical description of GMT. Figure 1 shows the framework of GMT, including a comparison between the one-off delta parameter drop strategy and the proposed GMT.
Figure 1: Illustration of our proposed method GMT, compared with the one-off drop delta parameters approach. The figure on the left delineates the distinction between a trivial drop and a random drop, wherein the trivial drop serves to diminish the parameter redundancy that arises during the fine-tuning process. Building upon this insight, we refine the training procedure by preferentially updating more significant parameters, as determined by the gradient information pertinent to the task-specific data. This selective updating is operationalized through the implementation of a masking strategy that filters out gradients with smaller absolute values.
2.1 Preliminary
Given a model parameterized by $\Theta=\{\theta_{ij}\}$, where $i$ and $j$ denote the index of a single parameter, we define a binary mask matrix $M=\{M_{ij}\in\{0,1\}\}$ with the same size as $\Theta$. Delta tuning focuses on updating an elaborately selected sparse subset of the model’s parameters, denoted as $\Theta_{\Delta}=\{\theta_{ij}\in\Theta|M_{ij}=1\}$, where $M_{ij}=1$ indicates selection for update and $M_{ij}=0$ indicates the parameter remains unchanged. The loss function $\mathcal{L}(\Theta)$ is computed based on the entire set of parameters $\Theta$.
In this way, the parameter update at iteration $t$ can be stated as:
$\theta_{ij}^{(t+1)}=\begin{cases}\theta_{ij}^{(t)}-\eta\cdot\nabla_{\theta_{ij}}\mathcal{L}(\Theta),&\text{if $M_{ij}=1$}\\ \theta_{ij}^{(t)},&\text{else}\end{cases}$ (1)
where $\eta$ denotes the learning rate, and $\nabla_{\theta_{ij}}\mathcal{L}(\Theta)$ is the gradient of the loss function with respect to the parameter $\theta_{ij}$. This update is selectively applied only to the parameters included in $\Theta_{\Delta}$, as indicated by the mask matrix $M$.
2.2 Gradient-Mask Tuning
To incorporate the information of task-specific data during training, we propose GMT to select important parameters, leading to a finer-grained level of parameter selection.
For each batch of training data, we compute the gradient of the loss function $\mathcal{L}(\Theta)$ with respect to the parameters. To facilitate gradient accumulation, we define an accumulation interval $N$, which denotes the number of mini-batches over which gradients are accumulated before a parameter update is performed. The accumulated gradient $\Gamma_{ij}$ for each parameter $\theta_{ij}$ is defined as:
$\Gamma_{ij}=\frac{1}{N}\sum_{n=1}^{N}\nabla_{\theta_{ij}}\mathcal{L}(\Theta,\mathcal{B}_{n})$ (2)
where $\mathcal{B}_{n}$ represents the $n$-th mini-batch of data and $\nabla_{\theta_{ij}}\mathcal{L}(\Theta,\mathcal{B}_{n})$ is the gradient of the loss function with respect to $\theta_{ij}$ for the mini-batch.
Upon accumulating the gradients over $N$ mini-batches, we employ gradient information as a signal to identify the importance of parameters at the element-wise level in a fine-grained manner. This operation involves sorting the components of each accumulated gradient $\Gamma_{ij}$ by their absolute values and selecting a predefined top percentile $k$ for updating the parameters. The masked gradient $\mathcal{M}(\Gamma_{ij},k)$ is calculated as:
$\mathcal{M}(\Gamma_{ij},k)=\left\{g_{ij}\ |\ g_{ij}\in\Gamma_{ij},\ |g_{ij}|\geq T_{k}\right\}$ (3)
where $g_{ij}$ represents the $ij$-th component of the accumulated gradient for parameter $\theta_{ij}$. The threshold $T_{k}$ is the value such that the absolute values of the components of $\Gamma_{ij}$ that are greater than or equal to $T_{k}$ fall within the top $k$ percentile of all components by magnitude.
The subsequent parameter update step utilizes the masked gradient $\mathcal{M}(\Gamma_{ij},k)$:
$\theta_{ij}^{(t+1)}=\theta_{ij}^{(t)}-\eta\cdot\mathcal{M}(\Gamma_{ij},k)$ (4)
where $\eta$ is the learning rate and $t$ indexes the current training step. The detailed algorithmic training procedure is given in Algorithm 1.
Algorithm 1 Gradient-Mask Tuning
Hyperparameters: The accumulation interval $N$, the training step $T$ and the learning rate $\eta$.
Input: The initial model parameters $\Theta^{(0)}$ and the $n$-th mini-batch of data $\mathcal{B}_{n}$.
Output: The updated model parameters $\Theta^{(T)}$.
1:for $t$ in $10\to T-$ 1 do
2:     $\Gamma\leftarrow 0$
3:     for $n$ in $1\to N$ do
4:         $\Gamma\leftarrow\Gamma+\nabla_{\Theta}\mathcal{L}(\Theta^{(t)},\mathcal{B}_{n})$
5:     end for
6:     $\Gamma\leftarrow\frac{1}{N}\Gamma$
7:     $T_{k}\leftarrow\text{Percentile threshold based on }\Gamma$
8:     $\mathcal{M}(\Gamma,k)\leftarrow\left\{g_{ij}\ |\ g_{ij}\in\Gamma_{ij},\ |g_{ij}|\geq T_{k}\right\}$
9:     for $\theta_{ij}\in\Theta$ do
10:         if $|\Gamma_{ij}|\geq T_{k}$ then
11:              $\theta^{(t+1)}_{ij}\leftarrow\theta^{(t)}_{ij}-\eta\cdot\Gamma_{ij}$
12:         end if
13:     end for
14:end for
Compared to the existing work on Parameter-Efficient Fine-Tuning (PEFT) (Xu et al., 2023; Han et al., 2024), which focuses on reducing computation and improving efficiency, GMT aims to identify important neurons in a model and elevate the upper limits of its performance through selective parameter updating.
3 Theoretical Analysis
We present the theoretical analysis in two parts. First, we explain why sparsity is feasible from an optimization perspective (Fu et al., 2023; Hui et al., 2024) by demonstrating that delta tuning with sparsity optimizes an upper bound on the original loss function in Appendix A. Then, we discuss the theoretical analysis of selecting the absolute value of the gradient as used in GMT.
In order to consider the impact of task-specific data on the importance of each parameter in a model, we identify and update only a sparse subset of parameters based on their impact on the loss function. We define the loss function $L(\Theta;\mathcal{D})$ over the dataset $\mathcal{D}$ as:
$\mathcal{L}(\Theta;\mathcal{D})=\frac{1}{n}\sum_{i=1}^{n}\ell(\Theta;(\mathbf{x}_{i},y_{i}))$ (5)
where $\ell$ represents the loss for a single data point, with $\mathbf{x}_{i}$ and $y_{i}$ denoting the input features and corresponding label, respectively. The full parameter set is denoted by $\Theta$.
To discern the impact of individual parameters $\theta_{ij}$ on the loss function $\mathcal{L}(\Theta;\mathcal{D})$, we consider the impact of their removal while keeping all other parameters constant. The differential effect of excluding a parameter $\theta_{ij}$ is quantified by the change in loss $\Delta\mathcal{L}_{ij}(\Theta;\mathcal{D})$, which is formulated as:
$\Delta\mathcal{L}_{ij}(\Theta;\mathcal{D})=\mathcal{L}(\mathbf{I}\odot\Theta;\mathcal{D})-\mathcal{L}((\mathbf{I}-\mathcal{E}_{ij})\odot\Theta;\mathcal{D})$ (6)
where $\mathbf{I}$ represents the identity matrix and $\mathcal{E}_{ij}$ denotes an indicator matrix that has the same dimensions as $\Theta$. In $\mathcal{E}_{ij}$, all elements are zero except for the $(i,j)$ element, which is one. The Hadamard product, indicated by $\odot$, performs an element-wise multiplication, isolating the effect of the single parameter $\theta_{ij}$.
Given the computational infeasibility of evaluating $\Delta\mathcal{L}_{ij}$ for each parameter, we invoke the first-order Taylor series expansion around the current parameter vector $\Theta$, which provides a linear approximation of the loss function’s behavior in the vicinity of $\Theta$. The first-order approximation is represented by the gradient of the loss function with respect to $\theta_{ij}$:
$\Delta\mathcal{L}_{ij}(\Theta;\mathcal{D})\approx\nabla_{\theta_{ij}}\mathcal{L}(\Theta;\mathcal{D})\cdot(-\theta_{ij})$ (7)
where $\nabla_{\theta_{ij}}\mathcal{L}(\Theta;\mathcal{D})$ denotes the partial derivative of the loss function concerning the parameter $\theta_{ij}$. The negative sign arises from the fact that we are considering the removal of the parameter, which corresponds to a negative perturbation in its value.
Our objective is to discern the important parameters within the network architecture that are relevant to task-specific data. To this end, we employ the magnitude of the gradient $\nabla_{\theta_{ij}}$ as the criterion for saliency. It is imperative to recognize that a high magnitude of the gradient (regardless of its sign) typically denotes that the parameter $\theta_{ij}$ exerts a substantial influence on the loss function, whether the effect is positive or negative. Consequently, such parameters must be preserved to facilitate learning on the corresponding weights. The absolute value of the gradient $\nabla_{\theta_{ij}}$ is then used as a saliency measure to determine the importance of the parameter:
$s_{ij}=\left|\nabla_{\theta_{ij}}\mathcal{L}(\Theta;\mathcal{D})\right|$ (8)
Using this saliency measure, we construct the binary mask matrix $M$ such that $M_{ij}=1$ if the parameter $\theta_{ij}$ is deemed important based on a predefined sparsity level $\kappa$:
$M_{ij}=\mathbb{I}[s_{ij}\geq\tilde{s}_{(\kappa)}]$ (9)
where $\tilde{s}_{(\kappa)}$ is the saliency threshold selected to ensure that only the top $\kappa$ influential parameters have their corresponding entries in $M$ set to one. The indicator function $\mathbb{I}[\cdot]$ yields one if the condition within the brackets is true and zero otherwise.
4 Experiments
Method HumanEval HumanEval+ MBPP MBPP+ Average GSM8k MATH
Mistral-7B Mistral-7B
Pre-trained 28.7 23.8 51.9 42.1 36.6 - -
SFT 68.3 64.0 56.1 46.9 58.8 75.3 27.0
One-off Drop 68.3 64.0 54.6 46.1 58.3 74.6 27.6
HFT 67.1 61.6 57.1 48.1 58.5 77.8 27.3
Random Mask 70.7 64.6 55.1 45.4 59.0 74.6 25.1
GMT 69.5(+1.2) 62.2(-1.8) 59.6(+3.5) 48.6(+1.7) 60.0(+1.2) 78.6(+3.3) 28.5(+1.5)
DeepSeek-Coder-Base-6.7B Llama3-8B
Pre-trained 47.6 39.6 72.0 58.7 54.5 - -
SFT 76.8 73.8 74.9 62.4 72.0 78.1 29.2
One-off Drop 76.2 72.6 74.7 62.4 71.5 78.1 29.4
HFT 74.4 70.1 75.2 62.9 70.7 74.2 28.5
Random Mask 72.6 67.1 78.8 67.5 71.5 80.3 31.3
GMT 78.0(+1.2) 74.4(+0.6) 75.7(+0.8) 63.7(+1.3) 73.0(+1.0) 82.0(+3.9) 32.0(+2.8)
Table 1: Experimental results for a single task in a specific domain. All models are evaluated with zero-shot prompting. We obtain the results of the code benchmark from EvalPlus  [^2] https://evalplus.github.io/leaderboard.htmland the results of the math benchmark from the corresponding original papers. “-” denotes that no zero-shot results were officially reported. Bold text indicates the best results for the fine-tuned model on each benchmark. An indication of the decrease and increase compared to vanilla SFT is shown in green and red, respectively.
4.1 Experimental Setup
We thoroughly evaluate the task-specific capabilities of models through various tasks in code generation, math reasoning and the general domain.
Training and Evaluation.
For code generation, we employ the Magicoder-Evol-Instruct-110K (Wei et al., 2023) as the training data, which is a decontaminated version of evol-codealpaca-v1333https://huggingface.co/datasets/theblackcat102/evol-codealpaca-v1. We utilize Mistral-7B  (Jiang et al., 2023) and DeepSeek-Coder-Base-6.7B (Guo et al., 2024) as the base models. The trained models are evaluated using the HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021) benchmarks, which are widely recognized for their effectiveness in measuring the proficiency of Python text-to-code generation. To enable a more comprehensive evaluation, we also introduce HumanEval+ and MBPP+, both of which are provided by EvalPlus (Liu et al., 2024b).
For math reasoning, the MetaMathQA  (Yu et al., 2023b) dataset is employed to fine-tune on the Mistral-7B and LLaMA3-8B [^4] https://github.com/meta-llama/llama3 models. The evaluation is conducted using the GSM8k (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021) benchmarks, which are specifically constructed to test the model’s capacity for mathematical reasoning and problem-solving.
For the general domain, the Tülu V2  (Wang et al., 2024b) dataset is utilized in training on the LLaMA2-7B  (Touvron et al., 2023) and LLaMA2-13B model. Following HFT (Hui et al., 2024), we evaluate model on MMLU (Hendrycks et al., 2020), GSM8k (Cobbe et al., 2021), BBH (Suzgun et al., 2023), TyDiQA (Clark et al., 2020), TruthfulQA (Lin et al., 2022) and HumanEval (Chen et al., 2021).
Implementation Details.
We choose different base models for different tasks. All training experiments were done on Nvidia A100 and Nvidia H100 machines. In addition, we utilize BFloat16 precision and set the weight decay to 0. We use the cosine learning rate scheduler after a linear warm-up stage with a ratio of 0.03.
Baselines.
In order to thoroughly evaluate the effectiveness of our method, we compare GMT to the following baselines:
  • SFT: Vanilla supervised fine-tuning.
  • One-off Drop: As an extension of (Yu et al., 2023a), dropping a preset ratio of trivial delta parameters after the vanilla supervised fine-tuning.
  • HFT [^5] The results of domain-specific HFT experiments for math reasoning and code generation are reproduced by us. Since it is not obvious which parameter selection strategy is better for HFT, we apply HFT in all parameters of the model using a category-level parameter selection strategy. (Hui et al., 2024): Half of the parameters are selected for learning the new task, and the other half is frozen to retain previous knowledge.
  • Random Mask: During training, a preset ratio of the delta parameters is randomly updated at a fine-grained element-wise level.
For fairness, we deploy the same mask ratio for One-off Drop, Random Mask and GMT, whose details are shown in Appendix B.
4.2 Main Results
Code Generation.
The experimental results for the code generation task are presented in Footnote 2. By integrating task-specific gradient information with avoidance of random parameter selection, GMT achieves an average performance improvement of 1.2% on the Mistral-7B model and 1% on the DeepSeek-Coder-Base-6.7B model. In contrast, HFT freezes half of the parameters to maintain the original capabilities of the model, and the learning process is hampered by the lack of sufficient parameter updates for a specific task domain. Furthermore, the random mask approach shows some benefits of using sparsity on the Mistral-7B. However, it fails to maintain this performance on another model, suggesting that it lacks robustness due to its inability to utilize task-specific data information. Compared to SFT, a One-off Drop directly on a fine-tuned model does not improve performance due to its sensitivity to the optimal drop ratio. In the highly professional field of code generation, this improvement highlights the efficacy of GMT in optimizing model performance through judiciously fine-grained selection of parameters.
Math Reasoning.
Footnote 2 presents a comparative analysis of the GMT method against various baseline approaches. GMT outperforms other methods by leveraging gradient information to capitalize on sparsity independent of the inconsistencies associated with random methods. Particularly, GMT performs excellently on the GSM8k benchmark, outperforming SFT by 3.3% and 3.9% on both models. In contrast, the HFT method shows an improvement over the baseline on Mistral-7B. However, it performs terribly on Llama3-8B, showing the instability of the method, which is attributed to the strategy of randomly chosen parameters. One-off Drop is slightly improved on the MATH benchmark but not as effective as the GMT method that performs the drop operation throughout the training process. In summary, this strategic use of gradient signal associated with the task-specific data ensures stable and progressive performance improvements throughout the learning process, as evidenced by its success across different models.
General Domain.
Table 2 shows the experimental comparison of the proposed GMT with the baseline. In the experiments with the fine-tuned Llama2-7B model, the GMT method outperforms the SFT by 3.0% on average across all tasks, and in particular, it significantly leads by 13.0% on GSM8k. GMT likewise demonstrated superior performance than HFT in multitasking scenarios in the general domain, obtaining an average lead of 1.1%. After the expansion of the LLM size, the GMT maintains its performance, with a 3.3% improvement in fine-tuning performance of Llama2-13B compared to the vanilla SFT. Notably, both Random Mask and HFT perform better than SFT, suggesting that the general domain of multi-tasking benefits from appropriate sparsity, even when the parameter selection strategy is completely random. With the same amount of parameter updates, GMT shows further improvement over Random Mask, suggesting that the strategy of selecting parameters based on gradient information derived from task-specific data in our method is practically effective.
Model Method MMLU GSM8k BBH TyDiQA TruthfulQA HumanEval Average
(0-shot, EM) (8-shot CoT, EM) (3-shot CoT, EM) (1-shot, F1) (0-shot, MC2) (0-shot, Pass@10)
Llama2-7B Pre-trained 41.6 12.0 39.9 48.4 38.5 26.2 34.4
SFT 48.5 25.0 42.2 51.2 41.7 36.9 41.0
HFT 50.8 30.5 43.6 52.3 45.4 34.6 42.9
Random Mask 47.4 34.5 44.4 52.9 47.9 33.7 43.4
GMT 47.6(-0.9) 38.0(+13.0) 43.3(+1.1) 53.1(+1.9) 47.5(+5.8) 34.6(-2.3) 44.0(+3.0)
Llama2-13B Pre-trained 52.2 34.5 50.7 50.3 49.8 32.7 45.0
SFT 50.6 45.0 47.8 55.0 42.6 42.4 47.2
HFT 54.5 46.5 53.7 56.7 45.7 43.5 50.1
Random Mask 54.6 50.5 52.8 56.5 45.2 41.4 50.2
GMT 54.6(+4.0) 54.0(+9.0) 51.5(+3.7) 57.1(+2.1) 46.0(+3.4) 39.5(-2.9) 50.5(+3.3)
Table 2: Results of Llama2-7B and Llama2-13B models finetuned on the Tülu V2 dataset. The results of Pre-trained, SFT, and HFT are taken from Hui et al. (2024). Bold text indicates the best results for the fine-tuned model on each benchmark. An indication of the decrease and increase compared to vanilla SFT is shown in green and red, respectively.
4.3 Further Discussions
Analysis of Mask Ratio.
Figure 2: Fine-tuning performance of the proposed GMT with respect to various mask ratios during training in three domains. The training LLMs utilized for code generation task, math reasoning task, and general domain are DeepSeek-Coder-Base-6.7B, Mistral-7B, and Llama2-7B, respectively. The experimental results for the MATH benchmark are presented on the secondary y-axis located on the right side of the figure.
To comprehensively ascertain the sensitivity of the hyperparameter mask ratio selection on GMT, we analyze the influence of the amount of parameter updates during training. Comparative experiments were conducted across three domains. We iterated through all experiments, applying the range of parameters to be fine-tuned with 10% granularity. For extreme cases, experiments were performed with mask ratios of 95% and 99%. As shown in Figure 2, the experimental results indicate that despite fluctuations in the number of parameter updates, the proposed method maintains a relatively consistent performance level with no more than 90% mask ratio. Our GMT reveals that the optimal performance in fields like mathematics and coding, which require specialized knowledge, is achieved with mask ratios of 20% to 40%. Even under extreme mask ratio settings, GMT maintains impressive performance in task-specific domains such as math reasoning and code generation. However, its effectiveness significantly diminishes in the general domain within multi-task scenarios. This suggests a stronger adaptability of GMT to single-task scenarios, where the entire learning process can be completed with merely a 1% parameter update. Conversely, in multi-task settings, the potential lack of sufficient parameter updates may lead to an inadequate learning process. Further analysis demonstrates that the performance of the proposed GMT exhibits robustness to variations in the mask ratio, despite the latter being an adjustable hyperparameter within the model’s configuration.
Analysis of Drop Strategy.
Figure 3: We employed three distinct strategies for selectively dropping delta parameters at various rates depending on the magnitude (absolute value) of the delta parameters: 1) preferentially dropping salient parameters, 2) preferentially dropping trivial parameters, and 3) dropping parameters randomly.
Our analysis reveals that not all delta parameters contribute equally to the model’s performance. Specifically, we find that parameters with salient values are critical for maintaining the efficacy of the LLM, and their removal leads to a notable and rapid degradation in model functionality. To systematically explore this phenomenon, we designed a series of experiments using the Mistral-7B model, focusing on domain-specific applications in mathematics (GSM8k and MATH). The experimental results are depicted in Figure 3.
Our results indicate substantial differences in the impact of each strategy on model performance. The strategy of dropping a portion of the delta parameters with small absolute values after vanilla fine-tuning demonstrates a relatively robust performance, only showing significant degradation when the dropout rate reaches 80%. This suggests potential benefits from sparsification, which may enhance performance by reducing overfitting or removing noise. In contrast, the random drop strategy leads to an apparent performance decline at a much lower sparsity level of 40%. The strategy of dropping parameters with salient magnitudes results in a precipitous decline in model capabilities. This finding underscores the importance of these significant delta parameters in sustaining the functional integrity of the model. Expectedly, the utilization of the gradient information to trivial drop during the training process can elevate the upper limit of LLM performance and is not constrained by drop rate.
Such experiments demonstrate that the parameter updates during the fine-tuning process of LLM are redundant, and that the use of a reasonable sparse updating strategy enables the model to learn better. Comparison with the experiments of random drop indicates that our strategy of considering task-specific data during training, using gradient information as a signal, and retaining neurons with large update magnitudes for completing the update is effective.
Model Method Traning Speed FLOPs
(samples/s) ($1e^{18}$)
Mistral-7B Code Generation Vanilla SFT 15.32 1.481
Random Mask 13.37 1.481
GMT 13.43 1.480
Llama2-13B General Domain Vanilla SFT 13.53 1.706
Random Mask 12.54 1.707
GMT 12.74 1.707
Table 3: Comparison of training speed and computational FLOPs, the number of train samples per second is an indicator of training speed.
Analysis of Efficiency.
To demonstrate the time-efficiency of proposed GMT, we compare the training speed and computed FLOPs [^6] All data logging is from https://wandb.ai. of GMT, vanilla SFT, and Random Mask, with the metric of number of training samples per second responding to the training speed, and the results are shown in Table 3. The FLOPs computed by the three tuning strategies during training are quite comparable due to the fact that the gradient information utilized by GMT is a by-product of model training and does not impose additional derivation operations. In terms of training speed, the random mask is close to GMT, being 14% and 6% slower than vanilla SFT in the code generation and general domain, respectively. Since GMT needs to calculate the threshold of the gradient being masked based on a preset ratio, a process that requires a TopK operation after the model backpropagates the gradient, and then drops values with smaller absolute values of the gradient based on the threshold. Nonetheless, this trade-off between a small minor time overhead and performance improvement is completely acceptable.
5 Related Works
Delta Tuning (DT) (Hu et al., 2022; Ding et al., 2023) stands out for its efficacy in tuning a subset of the LLM parameters, conserving the integrity of the predominant pre-trained weights. We categorized delta tuning into post-processing and in-training based on the stage of implementation. This section describes the properties of the typical methods of both strategies as well as the respective drawbacks.
Post-processing Strategy
The post-processing strategy, although requiring tuning with full parameter participation, achieves the same purpose as in-training through a series of delta parameter drop and reset strategies. Inspired by dropout, mixout (Lee et al., 2019) stochastically mixes source and target parameters as a regularization technique and extends to fine-tune downstream tasks. DARE (Yu et al., 2023a) finds that setting most (90% or even 99%) of the delta parameters to zero does not diminish the ability to fine-tune the LLM, and extends to model fusion accordingly. Ties-merging (Yadav et al., 2024) describes the delta parameter as a type of task vector that trims it, and then averages the parameters that are identical for multiple model aggregation symbols to achieve model merging. ExPO (Zheng et al., 2024) is inspired by model interpolation (Wortsman et al., 2022) and obtains better aligned models by direct extrapolation from the parameters of the aligned model obtained by DPO or RLHF and the initial SFT model. Although these methods can be implemented directly on fine-tuned models that are open access and require little additional computational overhead, they never achieve optimal performance due to the lack of refinement in the training process.
In-training Strategy
The in-training Strategy is to select a subset of parameters or additional extensions instead of full parameter fine-tuning. The Low-Rank Adaptation (LoRA) (Hu et al., 2021) leverages trainable rank-decomposition matrices to facilitate the fine-tuning process. Prefix-tuning (Li and Liang, 2021) adjusts the PLM by updating a set of prefixed parameters in every transformer layer that are closely related to the task, significantly reducing training costs and enhancing performance across multiple tasks. Adapter (Houlsby et al., 2019; Zhang et al., 2023; Hu et al., 2023) modules add small, trainable networks to each LLM layer, allowing specific fine-tuning while maintaining existing parameters’ stability. Moreover, methods such as Masking (Zhao et al., 2020), DiffPruning (Guo et al., 2021), BitFit (Zaken et al., 2022) and AdaLoRA (Zhang et al., ) have been proposed, which identify a subset of trainable parameters within the LLM to achieve efficient fine-tuning. (Pfeiffer et al., 2023) implements computational units typically as autonomous parameter-efficient modules. Information is routed to a subset of the modules and then aggregated.
These methodologies not only expedite the adaptation process of the model to new tasks but also substantially reduce the consumption of computational resources and training time. In contrast, our approach is focused on the task-specific data information and fine-grained tuning of parameters, which not only preserves the model’s generalizability but also significantly improves its adaptability and performance on specific tasks.
6 Conclusion
In this paper, we propose the Gradient-Mask Tuning (GMT), a pragmatic approach to delta tuning for LLMs. By leveraging gradient magnitudes associated with task-specific data to selectively update parameters, GMT strikes an optimal balance between maintaining computational simplicity and achieving fine-grained parameter tuning. Additionally, GMT identifies more significant parameters and eliminates the redundancy introduced by trivial delta parameters in the fine-tuning. We evaluate the performance of GMT on multiple models in the code generation, math reasoning, and general domain. Experiments demonstrate that GMT not only elevates the upper limits of LLM performance across diverse tasks but also exhibits robustness to various fine-tuning settings. We further analyze the computational efficiency of GMT to confirm that no additional computational FLOPs are needed, as well as acceptable extra time consumption. Moreover, GMT allows for delta tuning without destroying the network structure, and thus can readily replace SFT to accomplish the optimization process.
7 Limitations
This work has several limitations. First, the experiments in this paper could have tried more datasets and more base models. Furthermore, a well-performing LLM requires a sequence of training paradigms, and we have not yet considered the role of GMT in the pre-training and alignment training (e.g., via DPO or RLHF). Finally, this paper has not analyzed the effect of the amount of data on the results, i.e., the comparison of the results of GMT and SFT in the case of limited data.
Appendix A Effectiveness of Sparsity in Fine-tuning
We consider the task of minimizing the loss function $\mathcal{L}(\Theta)$ for a large language model (LLM) with parameters $\Theta$, subject to a sparsity constraint on the parameter updates. This can be formalized as the following optimization problem:
$\begin{gathered}\min_{\Theta}\mathcal{L}(\Theta)\\ \text{subject to}\quad\|(I-M)\Theta_{\Delta}\|^{2}=0,\end{gathered}$ (10)
where $\mathbf{I}$ represents the identity matrix, ensuring that the constraint applies across all dimensions of $\Theta$.
By introducing a vector of Lagrange multipliers $\lambda$, we construct the Lagrangian function $\mathcal{L}_{d}$ associated with our optimization problem (10):
$\mathcal{L}_{d}(\Theta,\lambda)=\mathcal{L}(\Theta)+\lambda^{\top}\|(\mathbf{I}-M)\Theta_{\Delta}\|^{2}.$ (11)
The dual problem seeks to find the saddle point of $\mathcal{L}_{d}$ through a min-max strategy:
¯$\bar{\mathcal{L}}=\max_{\lambda}\min_{\Theta}\mathcal{L}_{d}(\Theta,\lambda).$ (12)
The min-max inequality allows us to establish an upper bound for our primal problem. For the dual problem (12), the min-max inequality suggests:
$\min_{\Theta}\mathcal{L}(\Theta)+\|(\mathbf{I}-M)\Theta_{\Delta}\|^{2}$ (13)
$\leq\max_{\lambda}\min_{\Theta}\mathcal{L}_{d}(\Theta,\lambda)$
$\leq\min_{\Theta}\max_{\lambda}\mathcal{L}_{d}(\Theta,\lambda).$
Consequently, the right-hand side of inequality (13) provides an upper bound for the primal optimization problem:
$\min_{\Theta}\mathcal{L}(\Theta)\leq\max_{\lambda}\min_{\Theta}\mathcal{L}_{d}(\Theta,\lambda)=\bar{\mathcal{L}}.$ (14)
Given that the regularization term $\|(\mathbf{I}-M)\Theta_{\Delta}\|^{2}$ is non-negative, the dual formulation does not exceed the value of the original loss function $\mathcal{L}(\Theta)$, thereby establishing the desired upper bound. This ensures that the fine-tuning adheres to the imposed sparsity constraint, selectively updating parameters as dictated by the mask $M$, while maintaining the stability and integrity of the LLM throughout the adaptation process.
Appendix B Details of the training mask ratio
Table 4 describes the mask ratio details of the proposed GMT method for all models and all domain experiments. To be fair in comparison, One-off Drop, Random Mask and GMT utilize the same mask ratio. All experiments reach optimal performance at 10%-40% sparsity updates.
Model Domain Mask Ratio (%)
Mistral-7B Math Reasoning 20
Llama3-8B Math Reasoning 30
Mistral-7B Code Generation 10
DeepSeek-Coder-Base-6.7B Code Generation 30
Llama2-7B General 40
Llama2-13B General 10
Table 4: Details of mask ratio for all experiments in this paper.